OpenStudy (leahhhmorgannn):

You originally draw a design for an art contest on a 2 in. x 5 in. card. The second phase of the contest requires the drawing to be transferred to an 8.5 in x 11 in. standard sheet of paper and utilize as much of the space on the paper as possible. You determine that the largest size one of the dimensions of your drawing can be is 10.5 in. What is the length of the other dimension if the two drawings are similar? Type your exact answer in the blank without the units, and round to the nearest tenths.

OpenStudy (word2):

is this for flvs?

OpenStudy (leahhhmorgannn):

Geometry, yeah

OpenStudy (word2):

what segment?

OpenStudy (leahhhmorgannn):

1...

OpenStudy (word2):

if the 2" side of the card was scaled up to 10.5" then the scaling factor would be 10.5/2 = 5.25 . The 5" side would thus be then length of 5* 5.25 = 26.25 Which exceeds the maximum dimension of 11" , thus it has to be 5" side which is scaled up to 10.5 and the scaling factor therefore 10.5/5 = 2.1 . The scale up length of the 2" side is thus 2 * 2.1 = 4.2"

OpenStudy (leahhhmorgannn):

I was going to ask if it was 4.2 is that the answer?

OpenStudy (word2):

yeah that's the answer .

OpenStudy (leahhhmorgannn):

thanks.